# Is GPT-OSS Good? A Comprehensive Evaluation of OpenAI’s Latest Open Source Models thanks:

### v1

@@ IV-A Hardware Infrastructure
TABLE II: Overall performance summary across all evaluated benchmarks. Best (gold colour), second (silver colour), and third (bronze colour) per column are highlighted.
@@
IV-B Model Configuration
All models were evaluated in their original, unquantised form to ensure fair comparison and maximum performance. We specifically selected the strongest available open source model within a similar parameter range to provide a meaningful benchmark for GPT-OSS performance. The 70B parameter models (Llama 3.3 70B, DeepSeek-R1 70B) represent the most competitive alternative in the gpt-oss-120B’s size category, while smaller models such as Gemma 3 27B and Phi-4 Reasoning 14B provide context for the gpt-oss-20B variant. This selection ensures our comparison focuses on state-of-the-art open source models of comparable scale rather than an arbitrary baseline.
IV-B1 Generation Parameters
@@
IV-C Data Preparation
IV-C1 Dataset Loading and Preprocessing
Our evaluation employs ten diverse benchmarks to comprehensively assess model capabilities across multiple cognitive dimensions. Each dataset targets specific competencies, while collectively they provide a holistic view of model performance.
  • MMLU (Massive Multitask Language Understanding): 14,000+ questions spanning 57 subjects from humanities to STEM, testing broad knowledge and reasoning capabilities across academic disciplines. Stratified sampling ensures balanced representation across all subject areas.
  • GSM8K (Grade School Math): 8,500 linguistically diverse grade school mathematics word problems requiring multi-step reasoning and arithmetic operations. Chain-of-thought prompts were prepared to evaluate reasoning processes.
  • HumanEval: 164 hand-crafted Python programming challenges testing code generation, algorithmic thinking, and understanding of programming concepts. Complete test harness integration enables execution-based evaluation.
  • FinQA: Financial domain question-answering dataset with 8,000+ examples requiring numerical reasoning over financial documents and tables. Tests specialised domain knowledge and quantitative analysis capabilities.
  • PIQA (Physical Interaction QA): 20,000+ questions testing physical commonsense reasoning about everyday situations and object interactions. Evaluates understanding of physical world dynamics.
  • SciQ: 13,700 crowdsourced science exam questions covering physics, chemistry, and biology at middle and high school levels. Assesses scientific knowledge and reasoning abilities.
  • MedQA: 61,000+ medical examination questions from USMLE and other medical licensing exams across multiple countries. Tests specialised medical knowledge and clinical reasoning.
  • LegalQA: Legal reasoning benchmark with 20,000+ questions from bar examinations and legal case studies. Evaluates understanding of legal principles and analytical reasoning.
  • DialogSum: 13,460 dialogues with corresponding summaries, evaluating conversational understanding and summarisation capabilities. Tests the ability to extract key information from multi-turn conversations.
  • C-Eval: Comprehensive Chinese evaluation suite with 13,900+ questions across 52 disciplines, testing multilingual capabilities. Bilingual prompt templates ensure fair cross-lingual evaluation.
IV-C2 Data Quality Assurance
  • Manual inspection of 5% random sample from each dataset
  • Encoding verification for non-ASCII characters
  • Consistency checks for ground truth labels
  • Removal of corrupted or ambiguous examples
IV-D Execution Pipeline
IV-D1 Automated Evaluation Workflow
@@
IV-D2 Error Handling and Recovery
@@
IV-E Monitoring and Logging
IV-E1 Performance Monitoring
@@ IV-E2 Logging Infrastructure
IV-F Experimental Controls
IV-F1 Randomisation
Randomisation protocols minimise systematic bias in our evaluation. Models are evaluated in random order within each benchmark to prevent ordering effects, and test examples are shuffled to eliminate position-dependent biases. This approach ensures that observed performance differences reflect genuine model capabilities rather than experimental artefacts.
IV-F2 Blind Evaluation Protocol
To ensure the objectivity of our findings, particularly for benchmarks requiring manual quality checks, we implemented a blind evaluation protocol. This methodology is crucial for mitigating cognitive biases in human raters, such as confirmation bias or the expectancy effect, where prior knowledge of a model’s identity could influence assessment [83]. Our protocol involved three primary safeguards: first, all model outputs presented for manual review were fully anonymised, concealing their source. Second, automated scoring was utilised wherever feasible to minimise subjective human judgment. Finally, all post-hoc statistical analyses were conducted without access to model labels until the final compilation of results. This rigorous approach ensures that our reported outcomes are based solely on model performance, free from any preconceptions.
@@ V-A Performance Overview
Figure 3: Performance rankings across benchmark categories using general prompts. Error bars represent 95% confidence intervals following bootstrap methodology from Efron and Tibshirani [69]. Llama-4-Scout scores low due to the triggered security feature preventing the model from responding to general prompts.
Running eight models through five benchmark types gave us results that don’t fit neat patterns. Like Srivastava et al. [56] and Liang et al. [35] found, parameter count doesn’t predict performance. GPT-OSS breaks this rule spectacularly: the gpt-oss-120B model loses to its gpt-oss-20B sibling consistently, which shouldn’t happen according to scaling laws [3, 5].
Wei et al. [84] and McKenzie et al. [85] saw similar inverse scaling in other contexts, but not this broadly. Stats check per Dror et al. [64]: p << 0.01, Cohen’s d = 0.73, so it’s real (not noise). This messes with what Ganguli et al. [86] said about predictable scaling.
@@ V-B Knowledge and Understanding
MMLU is basically the go-to test for general knowledge [15, 56], and GPT-OSS doesn’t do great. Looking at different subjects, we see patterns like what Jiang et al. [87] found: MoE architectures have weird knowledge gaps.
Digging deeper, the knowledge gaps are predictable and Roberts et al. [88] saw this coming. GPT-OSS handles STEM okay (72% mathematics, 74% physics) but tanks on humanities and professional stuff. Same issue Kandpal et al. [89] found: what you train on is what you get.
@@ V-C Reasoning and Problem Solving
Mathematics problems via GSM8K [16] and extras from Hendrycks et al. [15] went better for GPT-OSS. Chain-of-thought prompting [58, 59] boosted scores by 15%, which matches what Wang et al. [90] found about reasoning tricks.
@@
Figure 4: Performance distribution across evaluation categories (collected from published benchmark results). Analysis methodology follows BIG-bench protocols [56] with visualization inspired by Burnell et al. [91].
The nature of errors provides insight into architectural limitations, supporting theories proposed by Dziri et al. [92] on reasoning failures in transformers. Both GPT-OSS models consistently fail on problems requiring numerical precision maintenance, particularly those involving unit conversions, echoing observations by Razeghi et al. [93] on numerical reasoning challenges.
@@ V-D Code Generation Capabilities
Code generation, evaluated through HumanEval [18] and MBPP [19], emerges as a relative strength for GPT-OSS models. This success aligns with findings from Xu et al. [94] and Chen et al. [18] on the relationship between model architecture and code generation capabilities.
The efficiency of GPT-OSS 20B in generating correct solutions with minimal tokens contradicts assumptions from Li et al. [20] about the relationship between model size and code conciseness. Analysis of error patterns reveals consistency with findings from Zhang et al. [95] on common failure modes in neural code generation, particularly in handling edge cases and complex data structures.
@@ V-E2 Quality Dimensions Analysis
Figure 5: Performance heatmap across model-benchmark combinations.
Figure 6: Detailed performance analysis of GPT-OSS models across task subcategories.
@@
V-E3 Key Findings and Implications
The evaluation revealed that while 75% of models successfully solved the logic puzzle, response quality varied dramatically. Notably, gpt-oss-120B demonstrated superior performance by recognising both valid solutions while maintaining excellent response quality with concise, well-structured output. This contrasts sharply with models featuring explicit chain-of-thought mechanisms (DeepSeek-R1, Phi-4 Reasoning), which exposed internal reasoning processes resulting in responses exceeding 26,000 characters, a 10-fold increase that severely impacts readability.
@@ V-E3 Key Findings and Implications
V-E4 Response Quality Distribution
Analysis of response characteristics revealed an optimal range of 2,000-3,000 characters for complex reasoning tasks, balancing completeness with human readability. Models within this range consistently achieved higher quality scores across dimensions of length appropriateness, readability, clarity, and conciseness. The GPT-OSS models’ performance in this task demonstrates their ability to produce human-friendly outputs without sacrificing correctness, a crucial factor for practical deployment often overlooked in traditional benchmarking.
This case study underscores a fundamental tension in current LLM design between reasoning transparency and usability. While exposing internal reasoning processes may aid in verification and debugging, our findings suggest that most practical applications benefit from what we term “selective transparency” and maintaining internal reasoning while presenting only refined conclusions to users. The GPT-OSS models’ balanced approach in this regard represents a pragmatic design choice that prioritises end-user experience without compromising solution quality.
@@ V-F Domain-Specific Performance
Evaluation across specialised domains reveals varied capabilities, as illustrated in Fig 5. Financial reasoning through FinQA shows GPT-OSS models achieving moderate performance, with the gpt-oss-120B at 84% accuracy. Scientific reasoning (SciQ), medical knowledge (MedQA), and legal understanding (LegalQA) benchmarks demonstrate consistent mid-tier performance. Physical reasoning through PIQA and conversational summarisation via DialogSum further confirms the models’ general-purpose nature without exceptional strength in any particular domain.
V-G Conversational Coherence
Multi-turn dialogue evaluation through MT-Bench [38] reveals moderate conversational capabilities for GPT-OSS models. The degradation pattern analysis, following methodologies from See et al. [96], reveals increasing confusion after 4-5 conversation turns.
V-H Computational Efficiency Analysis
@@
V-I Statistical Significance and Reliability
Rigorous statistical analysis following best practices from Dror et al. [64] and Card et al. [36] confirms the reliability of our findings. McNemar’s test [66] with Bonferroni correction [67] yields p-values below 0.05 for all reported differences. Effect sizes, calculated following Cohen [68] and interpreted according to Sawilowsky [100], range from medium (d = 0.52) to large (d = 1.84).
Bootstrap confidence intervals [69] constructed through 1000 iterations remain narrow, with a maximum width of ±2.1%, confirming stable estimates consistent with recommendations from Koehn [72]. Inter-rater reliability achieves $\kappa$ = 0.87 following Landis and Koch [101], indicating strong agreement. Sensitivity analysis following Ulmer et al. [102] shows minimal impact of hyperparameter variations on relative rankings.
V-J Efficiency Evaluation
[…]
V-K Mutlilingual Capability
The multilingual evaluation exposes significant limitations, with both GPT-OSS models achieving below 45% accuracy on Chinese-language tasks. Strong performance from models like Qwen 3, which incorporate language-specific optimisation, suggests that general-purpose training may be insufficient for true multilingual capability.
V-L Synthesis and Implications
@@
The most striking finding is the inconsistent scaling between gpt-oss-120B and gpt-oss-20B models. Contrary to established scaling laws, the smaller 20B variant outperforms its larger counterpart on multiple benchmarks, including MMLU (69% vs 66%) and SCIQ (87% vs 82%). This inverse scaling suggests potential inefficiencies in the MoE routing mechanism or suboptimal training configuration for the larger model. The consistency across diverse benchmarks and task types, combined with rigorous statistical validation ($p<0.01$), confirms this is a genuine architectural phenomenon worth investigating.
@@ VI Conclusion
We presented a comprehensive, multi-domain evaluation of OpenAI’s GPT-OSS models and analysed their behaviour within the broader open source landscape. Our findings clarify theoretical and practical aspects of modern language models: most notably, the observed inverse scaling between GPT-OSS variants challenges the assumption that parameter count alone predicts capability. Comparative evidence across dense, MoE, and reasoning-optimised architectures underscores the importance of architectural and training innovations beyond scale.
This study relies on established benchmarks that, while widely adopted, do not fully capture emerging capabilities or real-world robustness. We also did not exhaustively optimise prompting or decoding strategies for any single model. As the ecosystem evolves rapidly, results represent a snapshot in time. Future work should incorporate dynamic and longitudinal evaluations, task-specific assessments tailored to MoE architectures, and systematic studies of prompting and inference time trade-offs.
GPT-OSS models add architectural diversity and provide evidence to refine our understanding of MoE scaling dynamics. Insights from this evaluation can inform the design and deployment of sparse models and guide prioritisation of efficiency-aware research directions. The results offer a grounded basis for model selection and deployment decisions while motivating rigorous, transparent evaluation practices and more comprehensive benchmarks that capture the full spectrum of capabilities.

### v3

@@ IV-A Hardware Infrastructure
TABLE II: Overall performance summary across all evaluated benchmarks. Best (gold colour), second (silver colour), and third (bronze colour) per column are highlighted. We average our benchmark results with published results for a more moderate view.
@@
IV-B Data Preparation
To ensure reliable and consistent evaluation across all benchmarks, we implemented a comprehensive data quality assurance protocol prior to model evaluation. This multi-stage validation process was designed to identify and address potential issues that could compromise the integrity of our comparative analysis.
  • Manual inspection of representative samples: We conducted systematic manual review of a 5% stratified random sample from each dataset, totaling approximately 2,100 examples across all benchmarks. This inspection focused on verifying question clarity, answer correctness, and adherence to each dataset’s documented format specifications. Any identified inconsistencies or anomalies were flagged for further investigation and potential exclusion from the evaluation set.
  • Character encoding validation: Given the multilingual nature of several benchmarks, particularly C-Eval, we performed comprehensive encoding verification for all non-ASCII characters to prevent tokenization errors during model inference. This process involved UTF-8 encoding validation and detection of potential character corruption that could affect model performance, especially for Chinese language content and mathematical symbols present across multiple datasets.
  • Quality filtering and example removal: Based on our inspection findings, we systematically removed corrupted, ambiguous, or malformed examples that could introduce evaluation bias. This included questions with missing context, multiple valid answers, or unclear problem statements. The final evaluation set retained 98.7% of the original examples across all datasets, with removed examples documented to ensure reproducibility and transparency in our evaluation methodology.
IV-C Execution Pipeline
IV-C1 Automated Evaluation Workflow
@@
IV-C2 Error Handling and Recovery
@@
IV-D Monitoring and Logging
IV-D1 Performance Monitoring
@@ IV-D2 Logging Infrastructure
IV-E Quality Assurance
Reproducibility measures follow recommendations from Pineau et al. [pineau2021improving] and Dodge et al. [dodge2019show], including fixed random seeds, version locking, and comprehensive logging. Data contamination detection employs methods from Sainz et al. [sainz2023nlp] and Magar and Schwartz [magar2022data], while evaluation gaming prevention follows guidelines from Dekoninck et al. [dekoninck2024evading].
@@ V-A Performance Overview
Figure 3: Performance rankings across benchmark categories using general prompts. Error bars represent 95% confidence intervals following bootstrap methodology from Efron and Tibshirani [efron1994introduction]. Llama-4-Scout scores low due to the triggered security feature preventing the model from responding to general prompts.
Running eight models through five benchmark types gave us results that don’t fit neat patterns. Like Srivastava et al. [srivastava2022beyond] and Liang et al. [liang2022holistic] found, parameter count doesn’t predict performance. GPT-OSS breaks this rule spectacularly: the gpt-oss-120B model loses to or tie with its gpt-oss-20B sibling on nearly half of the benchmarks, which shouldn’t happen according to scaling laws [kaplan2020scaling, hoffmann2022training], as depicted in Fig 7.
Wei et al. [wei2023inverse] and McKenzie et al. [mckenzie2023inverse] saw similar inverse scaling in other contexts, but not this broadly. Stats check per Dror et al. [dror2018hitchhiker]: p << 0.01, Cohen’s d = 0.73, so it’s real (not noise). This messes with what Ganguli et al. [ganguli2022predictability] said about predictable scaling.
Fig 3 and Fig 4 presents a comprehensive view of model performance across all ten evaluation benchmarks, revealing distinct patterns in capability distribution. The visualization clearly demonstrates the heterogeneous nature of performance across different task domains, with no single model achieving consistent superiority across all categories. DeepSeek-R1 70B and Qwen 3 235B emerge as the strongest performers overall, achieving particularly notable results in mathematical reasoning (GSM8K: 96%) and conversational tasks (DialogSum: 94%). However, both GPT-OSS variants exhibit remarkably consistent mid-tier performance across most benchmarks, with the notable exception of their severe weakness in multilingual evaluation (C-Eval), where both models score below 30%. The inverse scaling phenomenon between GPT-OSS models is clearly visible across multiple benchmarks, with gpt-oss-20B consistently matching or exceeding gpt-oss-120B performance in MMLU (69% vs 64%) and SciQ (82% vs 87%). Interestingly, Llama 4 Scout shows dramatic performance variation, excelling in knowledge-based tasks like MMLU (99%) while failing in domain-specific evaluations such as DialogSum (0%), HumanEval (0%) and SciQ (0%), seemingly due to safety filtering mechanisms that impose restrictions to responses under certain contexts. The performance distribution underscores the importance of task-specific evaluation rather than relying on aggregate scores, as model rankings shift substantially across different cognitive demands.
@@ V-B Knowledge and Understanding
MMLU serves as the standard benchmark for general knowledge evaluation [hendrycks2021measuring], where GPT-OSS achieved suboptimal performance, ranking last with accuracy of 69% for gpt-oss-20B and 66% for gpt-oss-120B. Remarkably, the 20B model outperforms the 120B model, directly contradicting the conventional scaling laws and challenging assumptions about model capacity and performance relationships. This counterintuitive finding indicates that deeper investigation into the underlying mechanisms is needed.
Examination results over different subjects, reveals knowledge gaps. Diving deeper into the results, we found that the knowledge gaps are predictable: GPT-OSS demonstrates competent performance in STEM subjects (72% in mathematics, 74% in physics) while struggles with humanities and professional disciplines. This pattern aligns with findings from Kandpal et al. [kandpal2023large], reinforcing the principle that training data composition directly determines knowledge acquisition and retention.
@@ V-C Reasoning and Problem Solving
Mathematics reasoning via GSM8K [cobbe2021training] presents a contrasting picture, with GPT-OSS achieving accuracy of 82% for gpt-oss-20B and 88% for gpt-oss-120B. Unlike the inverse scaling observed in MMLU, GSM8K demonstrates conventional scaling behaviour, suggesting that scaling laws may be task-dependent rather than universal.
Chain-of-thought prompting [wei2022chain, kojima2022large] boosted scores by 15% for the 20B model, and 14% for 120B model, consistent with Wang et al.’s [wang2023self] findings on reasoning enhancement tricks.
@@
Figure 4: Performance distribution across evaluation categories (collected from published benchmark results). Analysis methodology follows BIG-bench protocols [srivastava2022beyond] with visualization inspired by Burnell et al. [burnell2023rethink].
Analysis of failure patterns reveals that both GPT-OSS models consistently fail on problems requiring numerical precision maintenance, particularly those involving unit conversions. echoing observations by Razeghi et al. These failures echo observations by Razeghi et al. [razeghi2022impact] regarding numerical reasoning challenges under zero-shot inference, potentially limiting practical applications requiring mathematical accuracy.
@@ V-D Code Generation Capabilities
Code generation, evaluated through HumanEval [chen2021evaluating], emerges as a relative strength for GPT-OSS models, achieving accuracy of 73% for gpt-oss-20B and 87% for gpt-oss-120B. This performance hierarchy aligns with findings from Xu et al. [xu2022systematic] on the relationship between model architecture and code generation capabilities.
The GPT-OSS 20B model demonstrates efficiency in generating correct solutions with minimal tokens, and analysis of error patterns reveals failure modes in neural code generation, particularly when handling edge cases and complex data structures.
@@ V-E2 Quality Dimensions Analysis
Figure 5: Performance heatmap across model-benchmark combinations. Darker blue indicates higher accuracy.
@@
V-E3 Response Quality Distribution
Analysis of response characteristics revealed distinct presentation strategies across models. The the GPT-OSS models which produced concise outputs in the 2,000-3,000 character range. In contrast, models explicitly designed for chain-of-thought reasoning (DeepSeek-R1, Phi-4 Reasoning) exposed comprehensive internal reasoning processes, resulting in responses exceeding 26,000 characters, a 10-fold increase that severely impacts readability.
It should be noted that this comparison reflects different design philosophies rather than inherent architectural advantages. Many reasoning-optimized models offer API parameters or prompting strategies to control thinking token visibility, suggesting that response length differences may represent presentation choices rather than fundamental capability differences. A fair comparison would require either accessing GPT-OSS internal reasoning processes or configuring all models to suppress detailed reasoning chains.
V-E4 Key Findings and Implications
The evaluation revealed that while 75% of models successfully solved the logic puzzle, response quality varied dramatically. Notably, gpt-oss-120B demonstrated superior performance over models featuring explicit chain-of-thought mechanisms (DeepSeek-R1, Phi-4 Reasoning), by recognising both valid solutions while maintaining excellent response quality with concise, well-structured output.
@@ V-F Domain-Specific Performance
Evaluation across specialised domains reveals varied capabilities, as illustrated in Fig 5 and Fig 7. Financial reasoning through FinQA shows divergent performance, with gpt-oss-120B achieving 84% accuracy versus gpt-oss-20B’s 58%. Legal understanding via LegalQA demonstrates similar conventional scaling (120B: 87% vs 20B: 70%), as well as physical reasoning through PIQA (120B: 78% vs 20B: 66%). However, DialogSum, MedQA and SciQA exhibit particularly noteworthy scaling anomalies. On DialogSum, both variants achieve identical performance (120B: 88% vs 20B: 88%), and medical knowledge assessment via MedQA shows equivalent performance (both models: 82%), suggesting a performance ceiling independent of model scale. More striking is the SciQA benchmark, where the smaller model outperforms the larger one (120B: 82% vs 20B: 87%), representing another clear violation of conventional scaling expectations. These findings further supporting the hypothesis that scaling laws may not universally apply across all cognitive domains.
These inverse scaling phenomena across multiple benchmarks (MMLU, SciQA) and comparable results (DialogSum, MedQA) collectively challenge the assumption that larger models consistently outperform smaller ones, highlighting the need for comprehensive investigation of scaling behaviors in large language models.
V-G Computational Efficiency Analysis
Fig 6 contextualizes efficiency by showing token count distributions across models. Reasoning-optimized architectures (DeepSeek-R1, Phi-4 Reasoning, Qwen3 235B) typically generate very long outputs (means 800–1,000 tokens, maxima above 1,600), while GPT-OSS models remain notably concise (means $\sim$200 tokens, medians below 100 for the 20B variant). Llama-family models occupy a middle ground. Such output-length differences directly influence compute cost, memory footprint, and readability, setting the stage for the efficiency results below.
@@ V-G Computational Efficiency Analysis
V-H Statistical Significance and Reliability
Rigorous statistical analysis following best practices from Dror et al. [dror2018hitchhiker] and Card et al. [card2020little] confirms the reliability of our findings. McNemar’s test [mcnemar1947note] with Bonferroni correction [bonferroni1936teoria] yields p-values below 0.05 for all reported differences. Effect sizes, calculated following Cohen [cohen1988statistical] and interpreted according to Sawilowsky [sawilowsky2009new], range from medium (d = 0.52) to large (d = 1.84).
Bootstrap confidence intervals [efron1994introduction] constructed through 1000 iterations remain narrow, with a maximum width of ±2.1%, confirming stable estimates consistent with recommendations from Koehn [koehn2004statistical]. Inter-rater reliability achieves $\kappa$ = 0.87 following Landis and Koch [landis1977measurement], indicating strong agreement. Sensitivity analysis following Ulmer et al. [ulmer2022deep] shows minimal impact of hyperparameter variations on relative rankings.
V-I Efficiency Evaluation
Memory and throughput. On A100/H100-class hardware, peak GPU memory stabilised around 80 GB per device for gpt-oss-120B (MoE; 8 experts active per token) versus only 16 GB for gpt-oss-20B under our batch/sequence settings (including 4-bit KV cache and activation checkpointing). This $5\times$ reduction enables higher-batch, lower-latency serving tiers and easier horizontal scaling. Throughput measurements confirm the efficiency gap: with identical decoding settings, gpt-oss-120B reached 128 tokens/s, while gpt-oss-20B sustained 178 tokens/s on matched H100 clusters.
@@ V-I Efficiency Evaluation
Design factors. Efficiency advantages derive from smaller active parameter counts (5.1B vs. 3.6B per token for 120B and 20B, respectively) and a reduced KV cache. Furthermore, as shown in Fig. 6, GPT-OSS models tend to produce competent outputs with less tokens, which turns into tangible cost savings at scale.
V-J Multilingual Capability
Multilingual evaluation highlights a major weakness of GPT-OSS: both variants scored poorly on Chinese-language benchmarks (C-Eval: 28% for gpt-oss-20B and 20% for gpt-oss-120B), well below the 45% threshold. In contrast, models incorporating language-specific optimisation, such as Qwen 3, achieve substantially higher accuracy. This gap indicates that general-purpose pretraining alone is insufficient for robust multilingual capability, particularly in non-English domains.
V-K Synthesis and Implications
Figure 7: Direct performance comparison between GPT-OSS variants across all evaluation benchmarks. Blue bars represent GPT-OSS-120B performance, orange bars represent GPT-OSS-20B performance.
@@ V-K Synthesis and Implications
The most striking finding is the inconsistent scaling between gpt-oss-120B and gpt-oss-20B models. Contrary to established scaling laws, the smaller 20B variant outperforms its larger counterpart on multiple benchmarks, including MMLU (69% vs 66%) and SCIQ (87% vs 82%), as depicted by Fig. 7 : the direct comparison of GPT-OSS variants. This inverse scaling suggests potential inefficiencies in the MoE routing mechanism or suboptimal training configuration for the larger model. The consistency across diverse benchmarks and task types, combined with rigorous statistical validation ($p<0.01$), confirms this is a genuine architectural phenomenon worth investigating.
@@ VI Conclusion
We presented a comprehensive, multi-domain evaluation of OpenAI’s GPT-OSS models and analysed their behaviour within the broader open source landscape. Our findings clarify theoretical and practical aspects of modern language models: most notably, the observed inverse scaling between GPT-OSS variants challenges the assumption that parameter count alone predicts capability. Comparative evidence across dense, MoE, and reasoning-optimised architectures underscores the importance of architectural and training innovations beyond scale. While this study relies on established benchmarks that, though widely adopted, do not fully capture emerging capabilities or real-world robustness, and we did not exhaustively optimise prompting or decoding strategies for any single model, the results represent a valuable snapshot in time of the rapidly evolving ecosystem. GPT-OSS models add architectural diversity and provide evidence to refine our understanding of MoE scaling dynamics, with insights from this evaluation informing the design and deployment of sparse models and guiding prioritisation of efficiency-aware research directions. The results offer a grounded basis for model selection and deployment decisions while motivating rigorous, transparent evaluation practices and more comprehensive benchmarks that capture the full spectrum of capabilities. Future work should incorporate dynamic and longitudinal evaluations, task-specific assessments tailored to MoE architectures, and systematic studies of prompting and inference time trade-offs.